OpenStudy (anonymous):

ason earned scores of 90 points, 82 points, 95 points, and 86 points on four tests. What is the lowest score he can get on the fifth test and still finish with an average score of 90 points? 95 points 96 points 97 points 98 points

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

Would you add these: 90 points, 82 points, 95 points, and 86 points

Directrix (directrix):

(Sum of first 4 + x) / 5 = 90

Directrix (directrix):

(Sum of first 4 + x) = 90 * 5 We will solve this for x when you get a sum for those four numbers. @sjman1